Compute the future value of an ordinary annuity with BD 110 semi-annually payments at 6% annual...

  1. Compute the future value of an ordinary annuity with BD 110 semi-annually payments at 6% annual interest for 23 years.

  1. Find the future value of an ordinary annuity with BD 750 quarterly payments at 8% annual interest for 9 years and 6 months.

  1. Determine the future value of an ordinary annuity with BD 50 monthly payments at 1% annual interest for 2345 days.

  1. On March 13, Nada joined a saving account. Her bank will automatically deduct BD 160 from her checking account at the end of each month, and deposit it into her saving account, where it will earn 5 % annual interest. The account comes to term on December 1. Find the following:

a. Find the future value of Nada’s saving account.

      b. Find Nada’s total contribution to the account.

c. Find the total interest earned on the account.

  1. Khalid recently set up a tax-deferred annuity to save for his retirement. He arranged to have BD 75 taken out of each of his monthly checks; it will earn 4% annual interest. He just had his 20 birthday, and his ordinary annuity comes to term when he is 60. Find the following:

a. Find the future value of Khalid’s annuity.

b. Find Khalid’s total contribution to the annuity.

c. Find the total interest earned on the annuity.

Marking Scheme:

5 Mark = Choosing the correct formula, Clear and Correct steps and final answer

4 Marks = Clear and Correct steps with wrong final answer

3 Marks = Choosing the correct formula with Clear steps

2 Marks = Choosing the correct formula with incomplete steps

1 Mark = Choosing the correct formula or Correct final answer

0 Mark = Wrong choosing of formula, wrong steps
Answer #1

i)

FV of Annuity = P*[{(1+i)^n}-1]/i

Where, P = Annuity = 110, i = Interest Rate = 0.06/2 = 0.03, n = Number of Periods = 23*2 = 46

Therefore, FV = 110*[{(1+0.03)^46}-1]/0.03 = 110*2.8950437/0.03 = $10615.16

Note: There are Multiple DIFFERENT Questions in the given question. And, As per GUIDELINES, we are supposed to Answer ONLY 1 QUESTION.